# FiDeLiS: Faithful Reasoning in Large Language Model for Knowledge Graph Question Answering

### v1

@@ Abstract
While large language models (LLMs) have achieved significant success in various applications, they often struggle with hallucinations, especially in scenarios that require deep and responsible reasoning. These issues could be partially mitigated by integrating external knowledge graphs (KG) in LLM reasoning. However, the method of their incorporation is still largely unexplored. In this paper, we propose a retrieval-exploration interactive method, FiDeLiS to handle intermediate steps of reasoning grounded by KGs. Specifically, we propose Path-RAG module for retrieving useful intermediate knowledge from KG for LLM reasoning. We incorporate the logic and common-sense reasoning of LLMs and the topological connectivity of KGs into the knowledge retrieval process, which provides more accurate retrieving performance. Furthermore, we propose to leverage deductive reasoning capabilities of LLMs as a better criterion to automatically guide the reasoning process in a stepwise and generalizable manner. Deductive verification serves as a precise indicator for when to cease further reasoning, thus avoiding misleading the chains of reasoning and unnecessary computation. Extensive experiments show that our method, as a training-free method with lower computational cost and better generality outperforms the existing strong baselines in three benchmarks. The code of this paper will be released later.
@@ 1 Introduction
The emergence and application of large language models (LLMs) [1, 31, 27] have attracted widespread attention from researchers and the general public. It demonstrates remarkable reasoning capabilities, managing to solve complex reasoning problems through step-by-step thinking and planning [47, 49]. However, the reasoning of LLMs is not invariably reliable and may conflict with factual reality [32], which limits the application in areas requiring high reliability, such as healthcare [14] and science [41].
@@
The knowledge graph (KG) stores high-quality domain-specific knowledge in structured triplets, such as Wikidata [43], YAGO [36], and NELL [2]. KGs offer structured, explicit, and updated factual knowledge, which provides a faithful knowledge source for reasoning. Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretive process transparent. Due to its reliability and interpretability, it is considered as a promising method to improve the reliability of LLM reasoning. Therefore, several attempts have been conducted to integrate KGs with language models [32, 29, 17, 38]. Among them, the knowledge graph question answering (KGQA) is a critical task to incorporate the knowledge of KG into reasoning models [12, 44, 55]. Concretely, KGQA faces the following main challenges:
(I) How to retrieve specific knowledge from KGs to help reasoning precisely? Existing solutions include direct retrieval [37, 21] and semantic parsing [39, 10]. Direct retrieval involves taking the questions as a query and the knowledge triplets in the KG as candidates, using either sparse or dense retrieval techniques to identify several candidates most relevant to the query. Semantic parsing, on the other hand, transforms the question into an executable structured query statement (e.g., SPARQL) and executes the query in KGs. However, in KGs, some triplet by itself lacks rich semantic information due to the difference of KG schema, such as some entities are represented as machine identifiers (MIDs), and direct retrieval make it difficult to understand how relevant it really is to the user query. This becomes especially tricky in multi-hop question answering, where the intermediate triplets seem not directly related to the query can actually be crucial links to finding the right answer. Furthermore, semantic parsing faces the issues of non-executable or incorrectly executed generated queries [55].
@@
(II) How to make the reasoning model understand and utilize the retrieved structured knowledge in KGs? Existing solutions include (1) finetune LMs on KG structure to generate relation path grounded by KGs. These reasoning paths are then used for LLMs to conduct faithful reasoning and generate interpretable results [29, 55]; (2) leverage the reasoning model to iteratively retrieve and reason on the subgraphs from KG [38, 9, 21] (i.e., at each step, the LLMs retrieve neighbors of the current nodes and then decide to answer the question or continue the next search step.) However, for the first approach, the reasoning steps generated by language models offer few guarantees about feasibility (i.e., the reasoning step might not exist in the KG), especially when multiple consecutive steps are combined into a reasoning path. We conduct the error analysis regarding this issue in Section 4. The latter solution face the challenge regarding determining the optimal stopping point for the exploration process. For example, ToG [38] prompts the LLMs to evaluate the adequacy of current reasoning paths for answer generation. However, assessing the adequacy of reasoning paths is itself a complex task that demands a deep understanding of the domain. This lack of clarity can easily result in either premature stopping or excessive continuation, further complicating the decision-making process in LLMs.
To address these two challenges, we propose a retrieval-exploration interactive method, FiDeLiS. It is designed to iteratively retrieve and reason on the subgraphs from KG, simulating a dynamic searching process. Specifically, we propose a Path-RAG module for retrieving useful intermediate knowledge from KGs. Compared to vanilla retriever which easily returns outputs that are not useful for finding the correct answer based on relevance to the query, we incorporate LLMs to generate an exhaustive list of keywords from the query to maximize coverage and ensure that no potential reasoning paths will be overlooked. We also integrate the topological connectivity of KGs to enhance system accuracy by considering entity relationships in the knowledge retrieval process. Furthermore, we propose to leverage deductive reasoning capabilities of LLMs [26, 18] as a better criterion for automatically directing the reasoning process step by step. Deductive reasoning serves as a clear indicator for ceasing further reasoning when necessary, which helps prevent errors in the reasoning chain and reduce unnecessary computation. Our framework enables us to employ beam search, a technique tailored for intermediate reasoning steps within a controllable searching space. We can guide the searching process by ensuring the logical soundness of each reasoning step to prevent errors throughout the reasoning process. Overall, our main contributions include:
  • We propose a retrieval-exploration interactive method that is specifically tailored to handle intermediate steps of reasoning grounded by KGs.
@@
  • We propose deductive verification as precise indicators for when to cease further reasoning, thus avoiding misleading the chains of reasoning and unnecessary computation.
  • We integrate the reasoning process and topological connectivity of KGs into the knowledge retrieval process, which aids in recalling useful intermediate knowledge for reasoning and maintains efficiency as KGs scale up.
@@
  • Extensive experiments show that our method, as a training-free method with lower computational cost and better generality outperforms the existing strong baselines in three benchmarks.
@@ 2 Method
Overview. FiDeLiS is designed as a retrieval-exploration interactive method as shown in Figure 1. It is designed to iteratively retrieve and reason on the subgraphs from KG, simulating a dynamic searching process through KG. FiDeLiS incorporates two modules Reasoning Path Retrieval-Augmented Generation (Path-RAG) and Deductive-Verification Guided Beam Search (DVBS), where Path-RAG is utilized to retrieve a highly pertinent set of nodes/relations from KGs to the user query and determine the reasoning step based on the relevance score discussed in Section 2.1. On the other hand, DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) for finding the most promising reasoning paths. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path extension. To prevent either premature stopping or excessive continuation of reasoning path extension, DVBS enhances the clarity of termination signals using deductive reasoning discussed in Section 2.2.
@@ 2.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
Path-RAG comprises three main steps: initialization, retrieval, reasoning step candidates construction, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
Initialization. We initiate the Path-RAG by generating node and edge embeddings using a pre-trained Language Model (LM). These embeddings are then stored in a nearest neighbor data structure. To elaborate, consider $e_{i}\in\mathcal{E},r_{i}\in\mathcal{R}$, where $\mathcal{E}$ refers to the entities and $\mathcal{R}$ represents the relations in the relational graph. We leverage the pre-trained LM (e.g., SentenceBert [33] and E5 [46]) to $e_{i}$ and $r_{i}$, yielding the representation $z(e^{i})$ and $z(r^{i})$:
$z(e^{i})=\operatorname{LM}\left(e_{i}\right)\in\mathbb{R}^{d},\\ z(r^{i})=\operatorname{LM}\left(r_{i}\right)\in\mathbb{R}^{d}$ (1)
where $d$ denotes the dimension of the output vector.
Retrieval. To retrieve the entities and relations that are relevant to the user query, the embeddings of them are then utilized to populate a nearest neighbor index. Specifically, given a query $q$, we leverage the LM to generate an exhaustive list of keywords or relation names $k_{i}\in\mathcal{K}$ that would be used to find the reasoning path from the knowledge graph to answer the query (as referred to prompt list in Section B). The generated list of keywords aims to maximize the coverage of potential reasoning steps and to enhance the subsequent decision-making process of LLMs by concentrating on a smaller set of relevant reasoning step candidates.
$\mathcal{K}=\text{LM}(\text{prompt}_{p},q)$ (2)
Then each $k_{i}^{m}\in\mathcal{K}$ is concatenated into a string $k_{i}$ and embedded into the same latent space by the LM to obtain $z(k)\in\mathbb{R}^{d}$, the retrieval process involves identifying the node or sequence of nodes that best aligns with the semantic content of the keywords as follow:
@@
where $z(e)$ and $z(r)$ are the embeddings of nodes $e$ and edges $r$, respectively. We use the cosine similarity function, $\cos(\cdot,\cdot)$, to measure the similarity between the keyword representation and the node/edge embeddings. The $\operatorname{argtopm}$ operation retrieves the top-m elements based on this similarity, providing a set of nodes $e_{k}$ and edges $r_{k}$ considered most relevant to the query. We denote the score of each entity and each relation as $S_{e}^{i}=\cos\left(z(k),z(e)\right)$ and $S_{r}^{i}=\cos\left(z(k),z(r)\right)$.
Reasoning Step Candidates Construction. To construct reasoning steps candidates based on the finite set of nodes $\mathcal{E}_{m}$ and edges $\mathcal{R}_{m}$ relevant to the query and the relevance score $S_{e}^{i}$ and $S_{r}^{i}$, we denote the reasoning step can be scored and retrieved based on the following criteria as follows. The goal of this scoring function is to leverage the triplets to represent the connectivity of the KGs. It provides some hints on whether the current operation/decision would lead to a malicious operation several steps later by incorporating the next-hop neighbor’s information.
$S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}\left(S_{r}^{j}+S_{e}^{j}\right)$ (4)
It should be noted that the entity and relation mentioned in the equation should be logically connected in the corresponding KG. $S(p)$ denotes the score of the reasoning step; $S_{r}^{i}$, $S_{e}^{i}$ denote the relevance score of the next step relation and the corresponding neighbor; $S_{r}^{j}$, $S_{e}^{j}$ denote the relevance score of one further hop of next step. Parameter $\alpha$ is used to balance between immediate accuracy and long-term prospects within the graph. A higher $\alpha$ value encourages the model to consider less immediate but potentially vital outcomes, While a lower $\alpha$ makes it focus primarily on the immediate next steps without considerable regard for the long-term impacts. The detailed description can be found in Appendix Algorithm 1.
Figure 1: Overview of FiDeLiS. The goal of FiDeLiS is to extend the multiple-step reasoning paths grounded by the external knowledge graph. We instantiate the $T$ step reasoning path extension as shown in the figure. First, Path-RAG is used to retrieve a highly pertinent set of nodes/relations from KGs to the user query and construct the reasoning step candidates. Then, the constructed reasoning steps candidates are forwarded to DVBS. DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates for finding the most promising reasoning step. Specifically, at timestep $T$, DVBS leverage the LLMs to choose the top-$k$ (here we set the $k=2$ for better clarification) reasoning steps and decide whether to continue the next search step or cease the reasoning path extension based on the deductive verification.
@@ 2.2 DVBS: Deductive-Verification Guided Beam Search
DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) for finding the most promising reasoning paths. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path extension. It comprises three main steps: planning, beam-search, deductive-verification, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
@@
Plan-and-Solve. Inspired by the recent works regarding planning capabilities of LLMs [35, 56], this step aims to provide more hints for subsequent LLM decision-making. Specifically, we prompt LLM to generate the planning steps for searching the most promising reasoning paths initially denoted as $p$. The detailed prompts can be found in Section B.
Beam Search. In multi-step reasoning, a reasoning path of $T$ steps is sequentially generated through several time steps as $R=\left[s^{1},s^{2},\ldots,s^{T}\right]=s^{1:T}$ as shown in Figure 1. However, different from the typical text decoding process where each step consists of a single token, we follow the [50] and view a sequence of pair $(r,e)$ as a single step, where $r$ is the relation and $e$ is the corresponding entity. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path. We present the beam search at each step is modeled as follows:
$\mathcal{H}_{t}=\text{Top}_{K}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},p):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\right\}$ (5)
@@
where $\mathcal{H}_{t-1}$ denote the reasoning path at the previous time step $t-1$; $s^{t}$ refers to the current reasoning step; $\mathcal{S}$ refers to all possible next steps $s^{t}$ candidates produced from the Path-RAG; $\text{LM}(s^{t}|q,s^{1:t-1},p)$ refers to the language models prediction for the next step $s^{t}$ given the current sequence $s^{1:t-1}$, the query $q$, and the planning context $p$; $\oplus$ refers to appending $s^{t}$ to the current sequence $h$; $\text{Top}_{K}$ refers to selecting the top $K$ (denoted as beam width in the experiment setting) sets of reasoning steps.
@@
Deductive Verification. To mitigate the reasoning path extension terminates at the wrong point, we propose to leverage deductive reasoning capabilities of LLMs [26, 18] as a better criterion $C(x,s^{t},s^{1:t-1})\in\left\{0,1\right\}$ to automatically guide the decision making of LLMs in Equation 5. Specially, we first leverage LLMs to convert the user query to a declarative statement (as shown in Appendix G.2) and then use LLMs to judge whether the declarative statement of query $x^{\prime}$ can be deduced from current reasoning step $s^{t}$ and the previous reasoning steps $s^{1:t-1}$. We utilize the same backend LLM with prompt referred in Section B and define the deductive-verification criteria as follows:
$C(q^{\prime},s^{t},s^{1:t-1})=\begin{cases}1&\text{if declarative statement ofquery }q^{\prime}\text{ can be deduced from }s^{t}\text{ and }s^{1:t-1},\\ 0&\text{otherwise.}\end{cases}$ (6)
The overall DVBS can be modeled as follows:
$\mathcal{H}_{t}=\text{Top}_{B}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},p):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\text{ and }C(q^{\prime},s^{t},s^{1:t-1})=1\right\}$ (7)
This criterion is an essential component of our method for leveraging LLMs to iteratively validate each step of reasoning to make sure it follows logically and matches the original query. It acts as a clear sign to stop further reasoning when necessary, which helps avoid mistakes in the reasoning process and reduces unnecessary computations.

### v2

# FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering

@@ Abstract
Large language models are often challenged by generating erroneous or ‘hallucinated’ responses, especially in complex reasoning tasks. To mitigate this, we propose a retrieval augmented reasoning method, FiDeLiS, which enhances knowledge graph question answering by anchoring responses to structured, verifiable reasoning paths. FiDeLiS uses a keyword-enhanced retrieval mechanism that fetches relevant entities and relations from a vector-based index of KGs to ensure high-recall retrieval. Once these entities and relations are retrieved, our method constructs candidate reasoning paths which are then refined using a stepwise beam search. This ensures that all the paths we create can be confidently linked back to KGs, ensuring they are accurate and reliable. A distinctive feature of our approach is its blend of natural language planning with beam search to optimize the selection of reasoning paths. Moreover, we redesign the way reasoning paths are scored by transforming this process into a deductive reasoning task, allowing the LLM to assess the validity of the paths through deductive reasoning rather than traditional logit-based scoring. This helps avoid misleading reasoning chains and reduces unnecessary computational demand. Extensive experiments demonstrate that our method, even as a training-free method which has lower computational costs and superior generality, outperforms established strong baselines across three datasets. The code of this paper will be released at https://anonymous.4open.science/r/FiDELIS-E7FC.
@@ 1 Introduction
The emergence and application of large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; LLaMa-v3, 2024) have attracted widespread attention from researchers and the general public. They demonstrate remarkable reasoning capabilities, managing to solve complex reasoning problems through step-by-step thinking and planning (Wang et al., 2022b; Wei et al., 2022b). However, the reasoning of LLMs is not always reliable and may conflict with factual reality (Pan et al., 2024), limiting their application in areas requiring high reliability, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
@@
Knowledge graphs (KGs) store high-quality knowledge in structured triplets, such as Wikidata (Vrandečić & Krötzsch, 2014), YAGO (Suchanek et al., 2007), and NELL (Carlson et al., 2010). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning. Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent. Due to its reliability and interpretability, it is considered as a promising method to improve the reliability of LLM reasoning. Therefore, several attempts have been conducted to integrate KGs with language models (Pan et al., 2024; Luo et al., 2024; Hu et al., 2023; Sun et al., 2023). Among them, knowledge graph question answering (KGQA) is a critical task to verify the effectiveness of incorporating the knowledge from KGs into reasoning models (He et al., 2021; Wang et al., 2023a; Yu et al., 2022). However, effectively integrating KGs into reasoning models for KGQA presents several challenges, including data sparsity, the complexity of query interpretation, and the need for advanced inference capabilities. More specifically, we focus on two main questions concerning the integration of KGs with reasoning models as follows:
(I) How to retrieve specific knowledge from KGs to allow more precise reasoning? Existing solutions include direct retrieval (Sun et al., 2019; Jiang et al., 2022) and semantic parsing (Sun et al., 2020; Gu & Su, 2022). Direct retrieval utilize the query to find relevant knowledge triplets within the KGs. However, these triplets sometimes lack comprehensive semantic information due to variations in KG schema, such as entities being represented as machine identifiers (MIDs), rather than descriptive labels. This lack of semantic richness can obscure the relevance of the retrieved information with the user query, and becomes especially tricky in multi-hop question answering, where seemingly unrelated intermediate triplets may be pivotal in deriving the correct answer. On the other hand, semantic parsing aim to transform user queries into executable structured queries (e.g., SPARQL), which are then run on KGs. However, these methods often grapple with issues related to the non-executability or incorrect execution of the generated queries (Yu et al., 2022), which undermines the relaibility of these methods.
@@
(II) How to make the reasoning model understand and utilize the retrieved structured knowledge in KGs? Existing solutions include (1) finetuning LMs to generate relation paths grounded by KGs. These reasoning paths are then used for LLMs to conduct faithful reasoning and generate interpretable results (Luo et al., 2024; Yu et al., 2022); (2) leverage the reasoning model to iteratively retrieve and reason over subgraphs from the KG, deciding at each step to either answer the question, or to continue the searching step (Sun et al., 2023; Gu et al., 2023; Jiang et al., 2022). However, for the first approach, the reasoning steps generated by language models are not guaranteed to exist in the KG, especially when multiple consecutive steps are combined into a reasoning path. We conduct error analysis in Section 3 and show that only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG. The latter solution faces the challenge of determining the optimal stopping point for the exploration process. For example, ToG (Sun et al., 2023) prompts LLMs to assess the adequacy of current reasoning paths for answer generation. However, assessing the adequacy of reasoning paths is itself a complex task that demands a deep understanding of the domain. This lack of clarity can easily result in premature stopping or excessive continuation, further complicating the decision-making process in LLMs.
To address these two challenges, we propose a retrieval augmented reasoning method, FiDeLiS. It is designed to enhance KGQA by anchoring responses to structured, verifiable reasoning paths. FiDeLiS is composed of two major components, Path-RAG, which retrieves chain of entities and relations from KGs in a effective manner (discussed in Section 2.1), and Deductive-verification Beam Search (DVBS), which conducts deductive-reasoning-based beam search to generate multiple reasoning paths leading to final answers (discussed in Section 2.2).
The Path-RAG module uses a keyword-enhanced retrieval mechanism that fetches relevant entities and relations from a vector-based index of KGs to ensure high-recall retrieval. Compared to a vanilla retriever which easily returns outputs that are not useful for finding the correct answer based on relevance to the query, we incorporate LLMs to generate an exhaustive list of keywords from the query to maximize coverage and ensure that no potential reasoning paths will be overlooked. Once these entities and relations are retrieved, the module constructs candidate reasoning steps which are then refined using a stepwise beam search discussed below.
Next, DVBS leverages the deductive reasoning capabilities of LLMs (Ling et al., 2024; Huang & Chang, 2023) as a clear-defined criterion for automatically directing the beam search process step by step to create the complete reasoning paths. A key characteristic of DVBS is its integration of natural language planning (Song et al., 2023; Zhou et al., 2022) and beam search. This combination enhances the process of choosing the best reasoning paths by providing additional hints for decision making in LLMs. In addition, we redesign how reasoning paths are evaluated by converting the scoring process into a deductive reasoning task. This allows the LLM to validate the reasoning paths through deductive verification rather than traditional logit-based scoring. This proposed deductive verification serves as presie indicators for when to cease further reasoning, thus avoiding misleading reasoning chains and reducing the computational resources required. Overall, our main contributions include:
  • We propose a retrieval augmented reasoning method, which enhances knowledge graph question answering by anchoring responses to structured, verifiable reasoning paths grounded by KGs.
  • We propose a step-wise keyword retrieval method that enhances the recall of relevant intermediate knowledge from KGs. This ensures that all the paths we create can be confidently linked back to KGs, ensuring that they are faithful and reliable.
@@
  • We propose deductive verification as precise indicators for when to cease further reasoning, thus avoiding misleading the chains of reasoning and reducing unnecessary computation required.
@@
  • Extensive experiments show that our method, as a training-free method with lower computational cost and better generality, outperforms existing strong baselines in three datasets.
@@ 2 Method
[…]
Overview. The framework of FiDeLiS is as shown in Figure 1. Given a query, it first retrieves the reasoning step candidates from KGs (as discussed in Section 2.1). To ensure high recall, our method first generates a list of keywords from the query. And then based on the generated keywords, it retrieves entities and relations from an index (where the entities and relations are also embedded by vector representation obtained from an LM). Next, following the retrieved entities, n-hop reasoning step candidates are created and scored based on semantics similarity and neighborhoods aggregation. Then, a deductive verification-based beam search (discussed in Section 2.2) is employed to generate the top-k reasoning paths leading to the answer. First, a natural language plan is generated from the query to provide more hints for LLMs decision making and then a beam search procedure is conducted to score each reasoning step. We redesign the scoring function by converting the process into a deductive reasoning task. This allows the LLM to validate connections between paths through simple verification rather than traditional logit-based scoring. This change helps prevent easily misleading reasoning chains and reduces the computational resources required.
Figure 1: Overview of FiDeLiS. The goal of FiDeLiS is to extend the multiple-step reasoning paths $T$ grounded by KG. The figure shows given a query, Path-RAG first retrieve entities and relations that are relevant to the query, and construct the reasoning path candidates. Then DVBS constructs reasoning paths by iterative extending these path candidates using beam search. Specifically, at timestep $t$, DVBS leverage LLMs to choose the top-$k$ (here we set $k=2$ as an example) reasoning steps and decide whether to continue the next search step or cease the reasoning path extension based on deductive verification. At last, FiDeLiS return the top-$k$ reasoning paths as the final output leading to the answers of the query.
@@ 2.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
Path-RAG aims to retrieve reasoning path candidates from KGs and comprises three main steps: initialization, retrieval, reasoning step candidates construction, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
Initialization. We initiate the Path-RAG by generating embeddings for entities (nodes) and relations (edges) using a pre-trained language model (LM). We begin by extracting entities ($e_{i}$) and relations ($r_{i}$) from the KG, denoted as $\mathcal{E}$ and $\mathcal{R}$, respectively. Each entity and relation is then encoded using a pre-trained language model such as SentenceBert (Reimers & Gurevych, 2019) or E5 (Wang et al., 2022a), yielding dense vector representations:
$z(e^{i})=\operatorname{LM}(e_{i})\in\mathbb{R}^{d},\ z(r^{i})=\operatorname{LM}(r_{i})\in\mathbb{R}^{d}$ (1)
where $d$ denotes the dimension of the output vector. These embeddings are stored in a nearest neighbor data structure, which facilitates quick retrieval of similar entities or relations based on their vector representations.
Retrieval. To retrieve the entities and relations that are relevant to the user query, the embeddings of them are utilized to populate a nearest neighbor index. Specifically, for a given query $q$, we use LLM to generate an exhaustive list of keywords or relation names referred to as $\mathcal{K}$ (the prompt is given in Section F). These keywords are generated to maximize the coverage of potential reasoning steps essential for answering the query. The list is derived as $\mathcal{K}=\text{LM}(\text{prompt}_{p},q)$. Then each keyword $\in\mathcal{K}$ is concatenated into a string denoted as $K$ and then encoded into the same latent space by the same LM used in initialization step to obtain $z(K)=\text{LM}(K)\in\mathbb{R}^{d}$.
To find the entities and relations that best match the keyword embeddings, we compute the cosine similarity $\cos(\cdot,\cdot)$ between $z(K)$ and the embeddings of entities ($z(e)$) and relations ($z(r)$) in the KG. We then select the top-$m$ elements that exhibit the highest similarity:
@@
Each entity and relation in the retrieved set is then assigned a score, reflective of its similarity to the keyword embedding as $S_{\text{ent}}(e)=\cos\left(z(K),z(e)\right)$ and $S_{\text{rel}}(r)=\cos\left(z(k),z(r)\right)$.
Reasoning Step Candidates Construction. After the initial retrieval process, which runs once at the start of the algorithm, we iteratively construct reasoning step candidates to extend the reasoning paths based based on entities $\mathcal{E}_{m}$ and relations $\mathcal{R}_{m}$ that are relevant to the given query. To guide the selection of these candidates, we propose another scoring function based on the derived relevance score $S_{\text{ent}}(e)$ and $S_{\text{rel}}(r)$. Our scoring function leverages triplets to represent the connectivity of KGs and assess whether extending a path in certain directions might lead to negative outcomes over multiple steps by incorporating next-hop neighbor’s information.
$S((r,e))=S_{\text{rel}}(r)+S_{\text{ent}}(e)+\alpha\max_{\forall(r_{j},e_{j})\in N(e)}\left(S_{\text{rel}}(r_{j})+S_{\text{ent}}(e_{j})\right)$ (3)
Here, $(r,e)$ refers to the reasoning step defined as a relation-entity pair. $S_{\text{rel}}(r_{j})$ and $S_{\text{ent}}(e_{j})$ are the similarity scores for relations and entities at the next step. The component $N(e)$ corresponds to entities reachable from $e$ within one hop. The factor $\alpha$ is used to balance short-term outcomes and long-term potential in reasoning paths: a higher $\alpha$ prioritizes paths with long-term benefits, even if they seem sub-optimal initially, while a lower $\alpha$ emphasizes immediate steps, potentially overlooking future impacts. Details of Path-RAG are provided in Algorithm 1.
@@ 2.2 DVBS: Deductive-Verification Guided Beam Search
DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) to find the top-$k$ most promising reasoning paths that lead to the answer to the user query. At each timestep, we leverage the LLMs to choose the top-$k$ reasoning steps and decide whether to continue extending the reasoning paths or cease the extension process. It comprises three main steps: planning, beam-search, deductive-verification, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
@@
Plan-and-Solve. Inspired by the recent works regarding planning capabilities of LLMs (Song et al., 2023; Zhou et al., 2022), we prompt LLM to generate the planning steps for answering the user query, denoted as $w$. This step aims to provide more hints for subsequent LLM decision making process. The detailed prompt can be found in Section F.
Beam Search. To enable multi-step reasoning, a reasoning path of $T$ steps is sequentially generated through several time steps as $\left[s^{1},s^{2},\ldots,s^{t},s^{T}\right]=s^{1:T}$, where $s^{t}$ represents the reasoning step at timestamp $t$. We constrain each reasoning step should within a set of candidates $\mathcal{S}^{t}$ to control the computation efficiency. We construct the reasoning step candidates $\mathcal{S}^{t}$ based on the scoring function defined in Eq. 3 by selecting the top-$m$ reasoning steps with the highest scores. At each timestamp $t$, we leverage the LLMs to select the top-$k$ reasoning steps from the reasoning step candidates $\mathcal{S}^{t}$. The beam search process at timestamp $t$ is modeled as follows:
$\mathcal{H}_{t}=\text{Top}_{k}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},w):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}^{t}\right\}$ (4)
@@
where $\mathcal{H}_{t-1}$ denote the reasoning path up to the previous timestamp $t-1$ and $s^{t}$ refers to the current reasoning step. $\mathcal{S}^{t}$ refers to the set of possible next-step candidates retrieved by the Path-RAG at timestamp $t$. $\text{LM}(s^{t}|q,s^{1:t-1},w)$ refers to the language models prediction for the next step $s^{t}$ given the previous sequence $s^{1:t-1}$, the query $q$, and the planning context $w$. The operator $\oplus$ appends $s^{t}$ to the current path $h$, and $\text{Top}_{k}$ selects the top-$k$ reasoning paths (with $k$ controlling the beam search width, where a larger $k$ typically yield better performance (see the analysis in Figure 2)).
@@
Deductive Verification. To terminate the reasoning path extension process at the right point, we propose to leverage the deductive reasoning capabilities of LLMs (Ling et al., 2024; Huang & Chang, 2023) as a criterion $C(x,s^{t},s^{1:t-1})\in\left\{0,1\right\}$ to automatically guide the decision making of LLMs in Equation 4. Specially, we first leverage LLMs to convert the user query to a declarative statement $q^{\prime}$ (as shown in Appendix F.6) and then use LLMs to judge whether it can be deduced from current reasoning step $s^{t}$ and the previous reasoning steps $s^{1:t-1}$. We utilize the same backend LLM with prompt referred in Section F and define the deductive-verification criteria as follows:
$C(q^{\prime},s^{t},s^{1:t-1})=\begin{cases}1,&\text{if }q^{\prime}\text{ can be deduced from }s^{t}\text{ and }s^{1:t-1},\\ 0,&\text{otherwise.}\end{cases}$ (5)
The overall goal of DVBS model can be represented as:
$\mathcal{H}_{t}=\text{Top}_{B}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},w):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\text{ and }C(q^{\prime},s^{t},s^{1:t-1})=1\right\}$ (6)
This criterion is an essential component of our method for leveraging LLMs to iteratively validate each step of reasoning, ensuring that each step logically follows from the preceding steps and aligns with the original query. The function effectively signals when to terminate further reasoning, enhancing accuracy and minimizing unnecessary computational efforts.